OpenStudy (anonymous):

help please...... What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 36°?

OpenStudy (anonymous):

If a single exterior angle measures 36, n = (360/36) = 10. Sum = (n-2) 180

OpenStudy (anonymous):

oh yea yea , i got this now thxs.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so is it 1440?